# Testing Non-Standard Neutrinos in Purely Leptonic Lepton Decays

### v1

@@ III Numerical Results and Discussions
We consider three sterile neutrino masses: 0.1, 0.5, and 1.0 GeV. The corresponding distributions of the parameters $\delta\Upsilon_{\ell 1}$, $\delta\Upsilon_{\ell 2}$, and $\delta\Upsilon_{\ell 3}$ as functions of $Y^{2}$ are shown in Fig. 2. These parameters include a common factor $\frac{1}{U^{2}_{\ell^{\prime}4}+U^{2}_{\ell 4}}$, whose magnitude can be constrained by experimental observations. This, in turn, restricts the potential influence of sterile neutrinos on the extended Pontecorvo-Maki-Nakagawa-Sakata matrix Maki:1962mu ; Pontecorvo:1967fh . We find that the distributions for $\delta\Upsilon_{\ell 2}$ and $\delta\Upsilon_{\ell 3}$ exhibit singularities. These arise because the corresponding SM values for $\Upsilon^{S}_{\ell 2}$ and $\Upsilon^{S}_{\ell 3}$ vanish, whereas their new physics counterparts do not. The singularity occurs near $Y^{2}\simeq m^{2}_{\ell^{\prime}}/2$. This indicates that this behavior manifests only for $m_{4\nu}^{2}<m^{2}_{\ell^{\prime}}/2$, a region deserving particular attention in experimental measurements. In other words, only sterlie neutrinos lighter than the paraent charged lepton could cause observable effects. In particular, due to the condition $m_{4\nu}^{2}<m^{2}_{\ell^{\prime}}/2$, the restriction from phase space makes leptonic decays of the $\tau$ a better candidate than those of the $\mu$ for probing sterile neutrinos.
@@ IV Summary
We have investigated the decays $\ell^{\prime-}\to\ell^{-}\bar{\nu_{\ell}}\nu_{\ell^{\prime}}$ and discussed their polarization-dependent asymmetries in order to probe the possible existence of sterile neutrinos. Our results show that sterlie neutrinos can cause singularities in the distribution of the proposed asymmetries parameters as function of $Y^{2}$.

### v2

@@ III Numerical Results and Discussions
We consider three sterile neutrino masses: 0.1, 0.5, and 1.0 GeV. The corresponding distributions of the parameters $\delta\Upsilon_{\ell 1}$, $\delta\Upsilon_{\ell 2}$, and $\delta\Upsilon_{\ell 3}$ as functions of $Y^{2}$ are shown in Fig. 2. These parameters include a common factor $\frac{1}{U^{2}_{\ell^{\prime}4}+U^{2}_{\ell 4}}$, whose magnitude can be constrained by experimental observations. This, in turn, restricts the potential influence of sterile neutrinos on the extended Pontecorvo-Maki-Nakagawa-Sakata matrix Maki:1962mu ; Pontecorvo:1967fh . We find that the distributions for $\delta\Upsilon_{\ell 2}$ and $\delta\Upsilon_{\ell 3}$ exhibit singularities. These arise because the corresponding SM values for $\Upsilon^{S}_{\ell 2}$ and $\Upsilon^{S}_{\ell 3}$ vanish, whereas their new physics counterparts do not. The singularity occurs near $Y^{2}\simeq m^{2}_{\ell^{\prime}}/2$. This indicates that this behavior manifests only for $m_{4\nu}^{2}<m^{2}_{\ell^{\prime}}/2$, a region deserving particular attention in experimental measurements. In other words, only sterile neutrinos lighter than the parent charged lepton could cause observable effects. In particular, due to the condition $m_{4\nu}^{2}<m^{2}_{\ell^{\prime}}/2$, the restriction from phase space makes leptonic decays of the $\tau$ a better candidate than those of the $\mu$ for probing sterile neutrinos.
Existing experiments have already collected an impressive number of $\tau$ lepton samples, for instance, Belle II has accumulated approximately $45\times 10^{9}$ $\tau$-pair events with only 8% non-$\tau$ background, while future colliders such as CEPC and FCC are projected to deliver $70\times 10^{9}$ and $170\times 10^{9}$ $\tau$ pairs, respectively, with even cleaner conditions yuan2021tau . These datasets present outstanding opportunities for precision studies of $\tau$ decays.
However, these samples are currently unpolarized. If future colliders were to incorporate polarized $e^{+}e^{-}$ beams, they could produce polarized $\tau$ leptons, whose polarization could be reconstructed from angular distributions relative to the beam axis. This would allow significantly more sensitive tests of the SM and deeper probes of new physics, including the sterile neutrino signatures proposed in this work. Our analysis therefore identifies a promising phenomenological target and motivates the development of polarized-beam capabilities at next-generation colliders.
@@ IV Summary
We have investigated the decays $\ell^{\prime-}\to\ell^{-}\bar{\nu_{\ell}}\nu_{\ell^{\prime}}$ and discussed their polarization-dependent asymmetries in order to probe the possible existence of sterile neutrinos. Our results show that sterile neutrinos can cause singularities in the distribution of the proposed asymmetries parameters as function of $Y^{2}$.